### v2

# DeepASA: An Object-Oriented Multi-Purpose Network for Auditory Scene Analysis

Abstract
We propose DeepASA, a multi-purpose model for auditory scene analysis that performs multi-input multi-output (MIMO) source separation, dereverberation, sound event detection (SED), audio classification, and direction-of-arrival estimation (DoAE) within a unified framework. DeepASA is designed for complex auditory scenes where multiple, often similar, sound sources overlap in time and move dynamically in space. To achieve robust and consistent inference across tasks, we introduce an object-oriented processing (OOP) strategy. This approach encapsulates diverse auditory features into object-centric representations and refines them through a chain-of-inference (CoI) mechanism. The pipeline comprises a dynamic temporal kernel-based feature extractor, a transformer-based aggregator, and an object separator that yields per-object features. These features feed into multiple task-specific decoders. Our object-centric representations naturally resolve the parameter association ambiguity inherent in traditional track-wise processing. However, early-stage object separation can lead to failure in downstream ASA tasks. To address this, we implement temporal coherence matching (TCM) within the chain-of-inference, enabling multi-task fusion and iterative refinement of object features using estimated auditory parameters. We evaluate DeepASA on representative spatial audio benchmark datasets, including ASA2, MC-FUSS, and STARSS23. Experimental results show that our model achieves state-of-the-art performance across all evaluated tasks, demonstrating its effectiveness in both source separation and auditory parameter estimation under diverse spatial auditory scenes. The demo video, samples and code are available at https://huggingface.co/spaces/donghoney22/DeepASA.
1 Introduction
Auditory scene analysis (ASA) seeks to extract information about individual sound sources from complex auditory environments bregman1984auditory . This includes identifying the class of each source, its onset and offset times, and its direction of arrival (DoA) brown1994computational . In humans, ASA is facilitated by the brain’s ability to organize sound into perceptual streams by integrating multiple auditory cues such as pitch, timing, and spatial location, in a complementary fashion cherry1953some ; bregman1994auditory ; SHAMMA2011114 .
Inspired by this ability, computational auditory scene analysis (CASA) research brown1994computational ; wang2006computational has aimed to replicate similar functionality using auditory cues like pitch, onset/offset, interaural level differences (ILDs), and interaural time differences (ITDs). With the rise of deep learning, a wide array of models has emerged to tackle specific ASA tasks, including audio tagging (AT) kong2017joint ; gong2021psla ; schmid2023efficient , sound event detection (SED) mesaros2021sound ; cakir2017convolutional ; shao2024fine , DoAE grumiaux2022survey ; diaz2020robust ; wangfn , and blind source separation (BSS) wang2018supervised ; luo2019conv ; kavalerov2019universal .
However, unlike the human auditory system, these task-specific models typically lack the capacity for relational reasoning across tasks and cues. As a result, they often fail when critical auditory cues are missing or degraded he2021deep . Recent work highlights the benefit of combining multiple auditory cues. For instance, target sound extraction (TSE), which uses class veluri2023real ; delcroix2022soundbeam , timestamp wang2022improving ; kim2024improving , and spatial cues jenrungrot2020cone ; gu2024rezero , outperforms universal sound separation (USS) in challenging scenarios kong2023universal ; zmolikova2023neural . Similarly, joint SED and DoAE in sound event localization and detection (SELD) has been shown to enhance performance by leveraging interdependent auditory information adavanne2018sound ; he2021sounddet .
Building on this evidence, we propose a general-purpose ASA model that emulates this cue integration process. Our model separates object-level auditory streams at an early stage and performs multiple downstream ASA tasks by exploiting the complementary relationships among estimated cues.
Contribution. We introduce DeepASA, a unified architecture for general auditory scene analysis, with two core contributions:
1. Object-Oriented Processing (OOP): We design an encoder-decoder framework that isolates features of individual sound objects. The encoder incorporates a dynamic short-time Fourier transform (STFT) module and a transformer-based feature aggregator. These are followed by an object feature separator that disentangles per-object auditory representations. These object-centric features are then passed to multiple sub-decoders, each responsible for estimating a specific auditory parameter (e.g., class, DoA, timestamp) without requiring manual association or permutation invariant training between estimated auditory parameters.
2. Chain-of-Inference (CoI) architecture: To emulate the human auditory system’s ability to recover missing cues by reasoning from others, we propose a chain-of-inference architecture. This component progressively refines object representations by fusing the outputs of ASA sub-decoders using a temporal coherence-based attention mechanism. It enables the model to reinforce incomplete or ambiguous estimates via complementary information.
We validate DeepASA on multiple spatial audio benchmarks. Ablation studies on the newly proposed ASA2 dataset [^1] https://huggingface.co/datasets/donghoney22/ASA2_dataset show that task-specific sub-decoders mutually enhance one another when conditioned on shared object features, yielding strong results with an SI-SDRi of 11.2 dB and a SELD score of 0.206. We also demonstrate that DeepASA pretrained by the ASA2 dataset can be generalized to various spatial audio benchmarks through fine-tuning with CoI. The fine-tuned DeepASA achieves state-of-the-art (SOTA) results, including an SI-SDRi of 18.5 dB on MC-FUSS for multichannel universal sound separation aizawa2023unsupervised ; wisdom2021s , a SELD score of 0.253 on STARSS23 shimada2023starss23 .
Figure 1: Comparison between (a) traditional track-wise processing and (b) proposed object-oriented processing
2 Method
2.1 Overall framework: DeepASA
The multichannel audio mixture $\mathbf{x}\in\mathbb{R}^{M\times N}$, where $M$ is the number of microphones and $N$ is the number of time samples of the waveform, can be modeled as the summation of $J$ reverberant foreground signals and background noise $\mathbf{v}$. The $j$-th foreground source can be further decomposed as the superposition of direct sound $\mathbf{s}_{j}$ and reverberation $\mathbf{h}_{j}$.
$\textbf{x}[n]$ $=\sum_{j=1}^{J}(\textbf{s}_{j}[n]+\textbf{h}_{j}[n])+\textbf{v}[n].$ (1)
We aim to estimate the auditory information of up to $J$ foreground sources, including classes, timestamps of onsets and offsets, DoA trajectories, and multichannel waveforms of the direct and reverb audio signals, as well as one multichannel noise signal, from the multichannel audio mixture.
Figure 2: Basic architecture of DeepASA
As illustrated in Figure 2, the overall architecture of DeepASA comprises three components: audio encoder, object separator, and sub-decoders. The audio encoder extracts basic features from the $M$-channel input mixture. The object separator separates audio features extracted from the audio encoder into $J+1$ object features, each of which encapsulates the auditory information of a single sound object. The sub-decoders then estimate the auditory information from each object feature.
The proposed framework separates each sound object at a feature level, which is referred to as object-oriented processing (OOP) in this work. One advantage of OOP is that the object-wise permutation of the separated features is consistently inherited across all sub-decoders. For example, the $j$-th object estimated by the audio, SED, DoA decoders all include the waveforms, SED, DoA information from the same sound source. The characteristic is the major difference to the conventional track-wise processing adavanne2018sound ; nguyen2022salsa ; shul2024cst ; Liu_CQUPT_task3a_report ; Yang_IACAS_task3a_report ; Du_NERCSLIP_task3_report , for which multiple source information can be aligned on the same track when they do not overlap in time. The OOP eliminates the requirement for pairing different auditory information and also enables the selective attention of estimated information across different objects and auditory information.
2.2 Audio encoder
Dynamic STFT with time-varying learnable window The audio encoder of DeepASA incorporates our proposed architecture, Dynamic STFT, for adaptive temporal focusing on the input waveform. Conventional STFT uses a fixed window at all times, yet the locations and durations of salient information within time frames can vary across the waveform. Accordingly, we employ a learnable Gaussian window whose mean $(\mu_{t})$ and standard deviation $(\sigma_{t})$ are predicted at every frame, forming $\boldsymbol{\mu}\in\mathbb{R}^{T}$ and $\boldsymbol{\sigma}\in\mathbb{R}^{T}$ vectors for the total number of time frames $T$. The frame-wise mean $(\mu_{t})$ aligns the center of the window on the most informative region, while $(\sigma_{t})$ sculpts its taper: a large $\sigma_{t}$ flattens the window toward a rectangle, sharpening spectral focus with a thin main lobe, whereas a small $\sigma_{t}$ contracts it toward an impulse, enhancing temporal focus.
Figure 3: Time-varying learnable window
The process of extracting $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$ is depicted in Figure 3. For each microphone channel, a parameter-shared 1D convolution with the same kernel size and stride as the window size $K$ and hop size $L$ used in STFT is applied, producing outputs $\mathbf{x}^{\prime}\in\mathbb{R}^{D\times T}$ for each microphone. Two parallel linear layers then project $\mathbf{x}^{\prime}$ to the mean and standard deviation, producing $\boldsymbol{\mu},\boldsymbol{\sigma}\in\mathbb{R}^{T}$ for every microphone channel.
Processing STFT with the time-varying learnable window, the multichannel waveform is converted into a complex spectrogram $\mathbf{X}\in\mathbb{R}^{2M\times T\times F}$ where $F$ is the number of frequency bins. These complex spectrograms are then fed directly into an Up-Conv block composed of a 2D convolution layer followed by layer normalization, to increase the channel dimension from $2M$ to $C$ while preserving the temporal and spectral dimensions.
2.3 Object separator
Feature aggregation As the feature aggregation block of the proposed network, we modified the DeFT-Mamba lee2025deft model designed for the USS task (Figure 4(a)). DeFT-Mamba is a comprehensive feature aggregation model utilizing Mamba and transformer layers to capture temporal, spectral, and interchannel relations in multidimensional data. To lighten the model, we adopt Mamba-FFN only for T-Hybrid Mamba and use a conventional feedforward network (FFN) for F-Hybrid Mamba. Then, we remove the unfolding process from the gated convolutional block (GCB) while keeping the gating module and the convolution kernel.
Object splitter The features arranged by the modified DeFT-Mamba are transformed into $J+1$ object features through the object splitting layer given by a 2D convolution kernel. The separated $J$ object features corresponding to foreground sources pass through the direct and reverb decoders of the audio decoder, the SED decoder, and the DoA decoder. The last object among the separated objects is treated as a noise object, which is separately utilized by the noise decoder in the audio decoder to estimate background noise.
2.4 Sub-decoders
The sub-decoders decode each object feature into the direct, reverberant, and noise audio signals, class, timestamp, and DoA outputs. The detailed architectures of sub-decoders are depicted in Figure 4(b).
Figure 4: Detailed architecture of (a) feature aggregation block, and (b) sub-decoders
MIMO Audio decoder The audio decoder aims to estimate a maximum of $J$ foreground source signals and 1 background noise signal corresponding to individual sound objects. In the conventional multi-input single-output (MISO) model, spatial information is lost in the decoded single-channel signal, and the audio decoder cannot assist DoAE in the DoA decoder. Therefore, the audio decoder in DeepASA is designed as a multi-input multi-output (MIMO) architecture, facilitating spatial information in the separated object features. To further assist the development of spatial information and dereverberation, the audio decoder is trained to separate direct sound $\textbf{s}_{j}$ and reverberant sound $\textbf{h}_{j}$ of each object across all microphone channels. In addition, for reducing the influence of background noise, the last object from the object separator is separately processed by the noise decoder trained to estimate the background noise v. When active foreground sources are fewer than $J$, inactive sources are trained to estimate zero target signals.
SED decoder The SED decoder decodes the object features into three different outputs related to the SELD task: a predicted class probability, a single-channel object activation (onset and offset) curve, and a sound event map presenting activation of individual classes across time. Ideally, a single object is only mapped to a single class label and activation curve, and the sound event map would be redundant. However, the object features are not perfectly separated in practice, so training with the auxiliary sound event map was useful for improving the SELD performance. During evaluation, only the predicted class probability and object activation curve are used to calculate the SELD scores. The SED decoder combines a pre-trained audio teacher-student transformer (ATST) shao2024fine with two separate convolutional recurrent neural network (CRNN) cakir2017convolutional -based branches summarizing time and frequency information. We employ the pre-trained ATST to utilize various features learned from AudioSet, a large-scale dataset containing a wide range of audio classes gemmeke2017audio . The features from ATST are combined with auxiliary features developed from another Conv2d layer. The concatenated features are analyzed in time by Time (T)-GRU and summarized to predict the sound event map. For predicting a single-channel activation curve, adaptive statistical pooling (ASP) okabe2018attentive and one fully connected layer are applied along the class dimension. The class probability is predicted by superposing the logit derived from Frequency (F)-CRNN, which prioritizes the spectral information using Conv2d layers and F-GRU, with the logit obtained from ASP. For inactive sources, the class decoder is trained to predict an equal probability of $1/C$ across all $C$ classes, while the sound event and timestamp decoders are trained to produce zero values for all time frames. The detailed specifications of the SED decoder are provided in Appendix A.
DoA decoder The DoA decoder has the CRNN structure and outputs a stream of DoA vectors in Cartesian coordinates ($x$, $y$, $z$). The DoA output estimates a DoA vector with a magnitude of 1 when a source is present, and assigns all $x$, $y$, $z$ values to 0 when no source is detected, allowing for the estimation of the activity and incident angle. DoA is estimated for each time frame, enabling the trajectory prediction for moving sources.
2.5 Chain-of-inference
Figure 5: Detailed architecture of chain-of-inference
The initial estimation through sub-decoders can include misalignment between auditory parameters. For example, the DoA vector stream can exhibit different onsets and offsets with the sound event map. These asynchronous estimations are refined from the object separation level through the fusion of initial estimations. This refinement step, chain-of-inference, consists of two processing steps: temporal coherence matching (TCM) and feature fusion (FF) (Figure 5).
Temporal coherence matching In TCM, temporal coherence between SED and DoAE is evaluated through the multi-clue attention mechanism. One branch of multi-clue attention takes SED as queries and DoA as keys and values, and the attention output is added to SED to refine the SED information. Conversely, the other multi-clue attention branch uses DoA as queries and SED as keys and values. The outputs of the two multi-clue attentions are passed through two individual linear layers to produce the $D$-channel clues and then superposed for fusion. The fused clue is interpolated by a transposed convolution layer to match its time dimension $T^{\prime}$ with that of object features ($T$).
Feature fusion In the feature fusion block, the fused clue is utilized to refine the object feature separation. TCM output is processed by a feature-wise linear modulation (FiLM) layer perez2018film , which outputs values of $\beta$ and $\gamma$ for each channel and time frame. The output of the FiLM layer is then combined with the output from the feature aggregation module ($\mathbf{z}$) to inject information required for object separation. The object features refined by the object splitter 2 are then plugged again into the sub-decoders 2 to improve ASA parameter estimation.
For efficient training of the CoI architecture, we utilize multi-stage training. That is, Net 1 of Figure 5 is trained first, then Net 2 is trained with Net 1 frozen.
3 Experimental Settings
3.1 Datasets
The proposed architecture is trained and evaluated on three spatial audio datasets: ASA2, MC-FUSS, and STARSS23. Auditory Scene Analysis dataset V2 (ASA2) provides ground truth signals and labels for all acoustic parameters addressed in this study, including comprehensive auditory information for moving source separation and scene analysis in noisy, reverberant environment. Compared to a previously proposed dataset lee2025deft , ASA2 provides a direct and reverberant audio signal, and the maximum number of foreground sources has increased from four to five. The balance in the number of audio clips has been changed to be proportional to the number of sources constituting the scene. Multichannel-Free Universal Sound Separation (MC-FUSS) aizawa2023unsupervised is a benchmark dataset for multichannel USS, including mixtures of 1–4 sound sources, with one being the background source. STARSS23 shimada2023starss23 is a dataset for the SELD task, addressing various sound events with 3.8 hours of audio-video data collected from real environments. This dataset generally requires data augmentation, so additional data augmentation of 20 hours was conducted using a spatial scaper roman2024spatial . All datasets were set to a 16 kHz sampling rate and a 4-second duration.
3.2 Training and model configuration
For the STFT, the Gaussian window of 40 ms length and hop size of 20 ms (50% overlap) was used. The initial weight for the mean and standard deviation of the window was set to 0 and 6.67 ms, ensuring $6\times\sigma$ covers the full window range. The number of feature aggregation modules was $6$, and the channel dimension for the DeepASA was $64$. The initial learning rate $4\times 10^{-4}$ was used for the Adam optimizer during pre-training, and $4\times 10^{-4}$ and $5\times 10^{-5}$ were used for fine-tuning to the MC-FUSS and STARSS23 datasets, respectively. The learning rate was halved if the validation loss did not decrease for five consecutive epochs. The model was trained for 100 epochs, and the model from the epoch with the lowest validation loss was selected as the final model. The batch size was set to 2. The training was performed on eight GeForce RTX 3090 GPUs.
The direct and reverb decoders in the audio decoder were trained with the SA-SDR loss von2022sa . We configured a reference channel among multichannel signals and applied different weights (1:0.1) to the losses of the reference and the other channels. The noise decoder was trained by the SI-SDR loss le2019sdr . The joint loss was calculated by summing the losses with weights 1:1:0.01 for direct, reverb, and noise decoders. The losses for the SED decoder were cross-entropy, binary cross-entropy, and mean-square-error (MSE), for the class decoder, timestamp decoder, and sound event decoder, respectively. The MSE loss was used for the DoA decoder, which was summed with all aforementioned losses with equal weights. In evaluation, SI-SDRi (dB) and SDRi (dB) were used for USS performance evaluation, while SELD metrics (error rate (ER,%), F1-score (F1,%), localization error (LE, degrees), localization recall (LR, %)) adavanne2018sound were employed for SED and DoAE. We measured the model complexity by the parameter size and computational complexity.
4 Results
4.1 Ablation study of DeepASA blocks
We conducted ablation studies to assess the contribution of each module in DeepASA. Each row block of Table 1 corresponds to the ablation of a specific component. For sequential consistency, the final configuration from the previous block was used as the baseline for the subsequent analysis.
Framework The first ablation study is about incorporating the MIMO separation task, SED, and DoA decoders. The baseline was the MISO USS model lee2025deft , and USS performance was measured based on the reconstruction of the reference channel. When MIMO USS is introduced, the separation performance of the reference channel decreases. This can be attributed to the features extracted for all channels rather than a single channel. Introducing the SELD task using the SELDNet adavanne2018sound sub-decoder enhances the USS performance as well, showing that auditory scene analysis can effectively guide the source separation task. The SELD performance is further improved when the CRNN architectures are employed as SED and DoA decoders. Notably, USS performance metrics are recovered to those of the baseline, indicating that a sophisticated SELD decoder can effectively improve SELD performance with minimal impact on the USS task.
Table 1: Ablation study on the proposed model using the ASA2 dataset. The models with (+) or (-) indicate the addition or removal of the corresponding blocks from the baseline. The (+) in the parameter size indicates the parameter size of the pre-trained ATST.
Model variation USS SED DoAE SELD $\downarrow$ Complexities
SI-SDRi $\uparrow$ SDRi $\uparrow$ ER $\downarrow$ F1 $\uparrow$ LE $\downarrow$ LR $\uparrow$ Param. MAC/s
Framework DeFT-Mamba-MISO lee2025deft 10.4 11.3 - - - - - 3.6 M 83.8 G
DeFT-Mamba-MIMO 10.0 10.9 - - - - - 3.6 M 83.8 G
(+) SELDNet adavanne2018sound 10.2 11.1 42.0 58.2 28.6 63.2 0.341 5.4 M 86.1 G
(+) SED, DoA decoder 10.4 11.4 40.0 60.3 22.9 65.7 0.317 7.2 M 88.4 G
Object
separator
(-) Unfold 10.3 11.4 39.8 60.5 22.2 66.1 0.314 7.2 M 88.0 G
(-) Unfold, F-Mamba 10.0 11.1 39.9 60.4 22.0 65.9 0.315 6.3 M 77.1 G
SED
decoder
(+) F-CRNN 10.3 11.2 38.0 61.1 21.9 68.8 0.301 8.1 M 90.7 G
(+) ATST + T-CRNN schmid2024improving 10.2 11.2 35.7 66.1 21.2 71.0 0.276 6.3 (+96.8) M 96.5 G
(+) ATST + T- &F-CRNNs 10.3 11.2 34.1 66.6 21.3 72.8 0.266 8.1 (+96.8) M 98.8 G
Audio
decoder
(+) Noise decoder 11.0 11.7 30.3 69.8 21.2 76.0 0.241 8.1 (+96.8) M 98.9 G
(+) Direct/reverb,
noise decoder
10.8 11.5 30.0 69.6 19.1 76.2 0.237 8.1 (+96.8) M 99.0 G
Dynamic
STFT
(+) Time-invariant window 10.7 11.4 30.2 69.8 19.3 76.0 0.238 8.1 (+96.8) M 99.0 G
(+) Time-variant window 11.0 11.7 28.8 70.2 18.5 76.9 0.230 8.2 (+96.8) M 99.1 G
[…]
Table 2: Ablation study for the chain-of-inference architecture
Model variation USS SED DoAE SELD $\downarrow$ Complexities
SI-SDRi $\uparrow$ SDRi $\uparrow$ ER $\downarrow$ F1 $\uparrow$ LE $\downarrow$ LR $\uparrow$ Param. MAC/s
without chain-of-inference 11.0 11.7 28.8 70.2 18.5 76.9 0.230 8.2 (+96.8) M 99.1 G
without
audio decoder 2
(+) SED branch 11.0 11.8 26.6 71.8 18.2 76.5 0.221 10.3 (+96.8) M 101.6 G
(+) DoA branch 11.0 11.8 28.2 70.6 17.6 76.0 0.228 10.3 (+96.8) M 101.6 G
(+) SED & DoA 11.0 11.7 26.5 73.0 17.3 77.2 0.214 12.1 (+96.8) M 104.0 G
[…]
Object separator Next, we examine the impact of downsizing on the baseline model (DeFT-Mamba-MIMO + SED, DoA decoder). The performance reduction by omitting unfold was negligible, showing only -0.1 dB difference in SI-SDRi. Furthermore, when F-Mamba is removed, a performance degradation of 0.3 dB in SI-SDR and 0.002 in SELD score is observed. However, this comes with significant lightweighting in parameter size (0.9 M) and computational complexity (10.9 GMac/s), so we adopted the conventional FFN in the F-HybridMamba block.
SED decoder The efficacy of more sophisticated sub-decoders is tested in this section. Recent work has shown that combining pre-trained ATST and CRNN achieves high performance in sound event detection schmid2024improving . The SELD score is also significantly improved when the baseline, combined with the ATST + T-CRNN decoder, is fine-tuned to the ASA2 dataset. This stresses the importance of using the class decoder specialized to the classification task. Similarly, we combined the F-CRNN architecture to capture and emphasize global frequency relations, improving the SELD score further up to 0.266. Since the improvement was less only with the F-CRNN decoder, we can conclude that capturing both local and global time-frequency relationships is critical for effective SED.
Audio decoder With the best SED decoder suggested above, we conducted the model analysis for the audio decoder. The first improvement was adding a noise decoder that estimates multichannel background noise. Both separation and SED performance are significantly improved when the noise decoder performing noise estimation is added. As analyzed in detail in Appendix B, the noise decoder enhances the separation performance of domestic sounds that share similar time-frequency characteristics with background noise. Next, separate decoders for the direct and reverberant sounds were employed to estimate $\mathbf{s}_{j}$ and $\mathbf{h}_{j}$, respectively, instead of $\mathbf{s}_{j}+\mathbf{h}_{j}$ estimated by the baseline audio decoder. Here, the USS performance was evaluated by measuring SI-SDRi and SDRi on $\mathbf{s}_{j}+\mathbf{h}_{j}$ for fair comparison to previous cases. While this change slightly reduces separation performance, DoAE performance (LE and LR) is markedly improved. This suggests that learning to separate direct audio positively impacts DoA accuracy by suppressing reverberant signals. A detailed analysis of the direct/reverb decoders is provided in Appendix C.
Dynamic STFT Lastly, the effectiveness of dynamic STFT was validated. The first comparison set is the model using a learnable but time-invariant window, which was designed by applying mean pooling over the temporal dimension before passing it through the linear layers of a time-varying learnable window. This model shows negligible performance improvement compared to the baseline STFT-based model. In contrast, the time-varying learnable window increases all performance metrics, demonstrating the importance of dynamic temporal adjustment of the window position and length in feature extraction and separation. The detailed comparisons with various dynamic convolution kernels are presented in Appendix D.
4.2 Ablation study of chain-of-inference (CoI)
The results of the CoI ablation study are presented in Table 2. The baseline model corresponds to the best-performing configuration without CoI, identified in Table 1. To isolate the effect of individual auditory cues, we ablated the CoI module by including only a single query branch, either the SED or DoA branch in the Temporal Coherence Matching (TCM) module. Additionally, the audio decoder 2 was removed from Net 2 to eliminate the influence of two-stage training on the audio separation task. In these ablated configurations, USS performance was evaluated using the output from audio decoder 1 only.
The results reveal that incorporating the SED branch improves sound event detection (as indicated by enhanced ER and F1-score), while incorporating the DoA branch improves DoA estimation (as shown by reduced localization error, LE). These findings suggest that cross-attention with SED or DoA queries selectively refines the corresponding auditory information. Notably, combining both branches (SED & DoA) leads to further improvements in both metrics, indicating that when one cue is unreliable, the other can compensate to support more accurate inference. In the final configuration, the full CoI architecture was restored, with Net 2 including audio decoder 2. In this setting, the USS output from audio decoder 2 was used for evaluation. The complete CoI architecture yields the best overall performance, achieving 12.0 dB SI-SDRi on the USS task and a SELD score of 0.206. These results confirm that the proposed CoI mechanism effectively enhances estimation across auditory domains by explicitly modeling interdependencies and enabling complementary cue integration. We run five trials and report the average and standard deviation for the final model. A more detailed analysis of CoI’s contribution is provided in Appendix E
4.3 Generalization to spatial audio benchmarks & comparison with SOTA models
Table 3: USS performance comparison on MC-FUSS dataset
Model J=2 J=3 J=4 Total Param. MAC/s
ByteDance-uss kong2023universal 14.8 14.4 12.7 14.0 28.0 (+80.7) M 40.1 G
MC-BSRNN gu2024rezero 15.7 15.2 11.4 14.1 12.2 M 15.3 G
TF-GridNet wang2023tf 17.2 16.1 12.5 15.3 14.7 M 462 G
DeFTAN-II lee2024deftan 17.6 16.3 12.8 15.6 4.1 M 66.1 G
SpatialNet quan2024spatialnet 17.8 16.5 13.1 15.8 7.3 M 71.8 G
DeFT-Mamba lee2025deft 18.4 17.1 13.8 16.4 4.2 M 58.2 G
[…]
Table 4: SELD performance comparison on STARSS23 dataset
Model ER $\downarrow$ F1 $\uparrow$ LE $\downarrow$ LR $\uparrow$ SELD $\downarrow$
CST-former shul2024cst 59.0 42.6 20.5 61.3 0.416
MFF-EINV2 mu2024mff 54.0 42.5 18.7 62.6 0.398
CST-former2 shul2025cst 42.0 59.7 15.6 68.4 0.301
EINV-2 w/ data augmentation chain (2nd rank) Liu_CQUPT_task3a_report 42.0 57.5 15.8 72.7 0.301
NERC-SLIP System (1st rank) Du_NERCSLIP_task3_report (w/o Ensemble)
40.0 64.0 13.4 74.0 0.277
NERC-SLIP System (1st rank) Du_NERCSLIP_task3_report (w/ Ensemble)
38.0 66.0 12.8 75.0 0.260
[…]
We evaluated the generalization capability of DeepASA on two spatial audio benchmarks: MC-FUSS for USS and STARSS23 for SELD, and compared it to SOTA models. DeepASA was pre-trained on the ASA2 dataset and fine-tuned using the ground-truth parameters available in each target dataset. Since no existing model can simultaneously perform USS, SED, and DoAE tasks, comparison models were trained individually for the subset of tasks they support.
USS performance comparison on MC-FUSS dataset Table 3 presents a comparison between DeepASA (excluding the CoI, SED, and DoA decoders) and existing SOTA models for universal sound separation (USS). Despite being unable to utilize the CoI module during fine-tuning on MC-FUSS, DeepASA outperforms existing USS models, particularly in scenes with a larger number of foreground sources. This performance gain is largely attributed to the noise decoder’s explicit estimation of background noise and the benefits of fine-tuning from pre-trained weights.
SELD performance comparison on STARSS23 dataset We evaluate two versions of DeepASA (with and without CoI) against SOTA SELD models, including top-ranking submissions from the DCASE 2023 challenge. DeepASA achieves the highest overall performance on the STARSS23 dataset, even without employing the audio decoder. Both versions of DeepASA surpass the performance of the top-ranked model Du_NERCSLIP_task3_report , which relied on class-dependent separation networks and model ensembling. It is worth noting that DeepASA was pre-trained on the ASA2 dataset prior to fine-tuning. To control for the effect of pre-training, we further compared DeepASA to publicly available models that were also trained using ASA2. As detailed in Appendix F, DeepASA consistently achieves SOTA performance across all evaluated tasks under the same training conditions. These results collectively demonstrate the strength of DeepASA across diverse ASA tasks and datasets, validating the effectiveness of combining class, timestamp, and DoA trajectory information through the object-oriented processing (OOP) and chain-of-inference (CoI) architecture.
5 Related Works
Universal Sound Separation (USS) USS addresses the challenge of separating various sound sources in an auditory scene where multiple sources overlap kavalerov2019universal ; wisdom2021s . Humans perform source separation by leveraging various characteristics of individual sources, such as class, timestamp, and DoA. However, existing USS methods rely on time-frequency representations, which may limit their ability to exploit such source-specific attributes kong2023universal ; gu2024rezero . To address this, the proposed method draws inspiration from human segregation processing and integrates multiple audio cues within a unified framework, resulting in improved performance on the USS task.
Target Sound Extraction (TSE) TSE refers to the separation of specific sound sources from audio mixture, with a focus on the auditory information of the source veluri2023real ; delcroix2022soundbeam . TSE plays a significant role in the source separation area, where spatial and temporal characteristics are leveraged to extract the target sound wang2022improving ; kong2023universal . A limitation of TSE is that humans must explicitly provide clues such as class, timestamp, and DoA. However, the network itself needs to iteratively infer these clues and use them to refine the separation process. To overcome this limitation, the proposed method introduces a strategy in which the clues inferred by the network are integrated to improve separation performance.
Sound Event Localization and Detection (SELD) The SELD task involves performing both SED and DoAE within a single model adavanne2018sound ; nguyen2022salsa ; shul2025cst . Nevertheless, most previous SELD models first analyze auditory cues using separate DNN branches for SED and DoA, and then perform track-wise separation into auditory stream shimada2021ensemble ; cao2021improved ; vo2024resnet ; mu2024mff ; mu2024seld . This approach can lead to incorrect pairing of SED and DoA outputs, and especially causes performance degradation when handling polyphonic audio. Additionally, models like Du_NERCSLIP_task3_report have improved performance by incorporating a pre-trained target sound extraction model luo2019conv but cannot reuse features learned from separation for SELD. Instead, the proposed architecture reutilizes multiple auditory information estimated from common per-object features to refine the separation and SELD performance.
6 Conclusion
We proposed DeepASA for unified auditory scene analysis and separation, employing a multi-decoder architecture that utilizes relations between auditory information for both source separation and analysis tasks. The model leverages chain-of-inference to refine performance in USS, SED, and DoAE tasks. Experimental results demonstrated the effectiveness of sub-decoders and multi-clue attention in improving feature extraction. Our results show significant performance improvements over existing methods in various downstream datasets.
Broader impacts This paper focuses on the framework that emulates how humans perform ASA by utilizing various types of auditory information to segregate sound objects to solve the cocktail party problem. We believe that DeepASA offers a unified framework that not only advances the state of the art in auditory scene analysis and sound separation but also opens new directions for research at the intersection of spatial hearing, multi-task learning, and object-centric audio modeling.
Limitations and future work The main limitation is the large parameter size of the ATST adopted for the SED decoder. In this work, the separated object features are reprocessed by the patch-based ATST to keep the advantage of ATST pre-trained on a large set of audio classification datasets. Future research could explore pre-training the novel classification layer designed to be compatible with object features, using a large audio dataset like ATST.
Acknowledgements and Disclosure of Funding
This work was supported by the National Research Foundation of Korea (NRF) grant (No. RS-2024-00337945) and STEAM research grant (No. RS-2024-00464269) funded by the Ministry of Science and ICT of Korea government (MSIT), the BK21 FOUR program through the NRF grant funded by the Ministry of Education of Korea government (MOE).
Appendix / Technical Appendices and Supplementary Material
This appendix is organized as follows:
  • Appendix A provides detailed specifications for the SED decoder.
  • Appendix B provides an in-depth analysis of the impact of the noise decoder on classification.
  • Appendix C presents experiments on the effects of the direct and reverb decoders on DoAE.
  • Appendix D examines the experimental results using the time-varying learnable window.
  • Appendix E presents an analysis of the chain-of-inference mechanism.
  • Appendix F compares the proposed method with SOTA models on the ASA2 dataset and includes a real-world demonstration.
Appendix A Detailed specifications of the SED decoder
In this section, we provide detailed specifications of the SED decoder. The SED decoder consists of three branches: ATST, T-CRNN, and F-CRNN. The specifications of the SED decoder are presented in Table 5. First, in the ATST branch, the object features are converted into patches of length $T_{p}=4,F_{p}=64$. The sequence length corresponds to the number of these transformed patches, and the patches are reshaped along the channel dimension. Since the pre-trained ATST takes a single-channel mel spectrogram as input, a linear layer is added before ATST to match the embedding dimension, followed by adaptive pooling to align the time frame to $T^{\prime}=40$. In the T-CRNN branch, seven convolution layers are used, as in [45], followed by a linear layer after concatenating the output of the ATST branch. The channel dimension and pooled frequency dimension are then reshaped into a single dimension, with the time dimension treated as the sequence dimension, and passed through a T-GRU. Finally, in the F-CRNN branch, pooling is focused on the time dimension, followed by two convolution layers, after which the channel dimension and pooled time dimension are reshaped into a single dimension, and the frequency dimension is treated as the sequence dimension before passing through an F-GRU. The combined output of the ATST and T-CRNN branches then passes through a fully connected layer with an output dimension of 1 to produce the timestamp output, a fully connected layer with an output dimension of $C=13$ (the number of classes) to produce the sound event map, and is subsequently combined with the output of the F-CRNN after passing through attentive statistics pooling (ASP) [40] and a fully connected layer, producing the final class output.
Table 5: Detailed specifications of the SED decoder
SED decoder
ATST branch T-CRNN branch F-CRNN branch
patchification
$\frac{TF}{T_{p}F_{p}}\times DT_{p}F_{p}$
(3,3)@64, LN, Mish (3,3)@64, LN, Mish
Pool 2×5
(3,3)@64, LN, Mish
Pool 2×1 Pool 4×5
(3,3)@64, LN, Mish
FC 768 Pool 2×1
Pre-trained ATST (3,3)@64, LN, Mish (3,3)@64, LN, Mish
Pool 2×1
(3,3)@64, LN, Mish
AdaptivePool1D(250 → 40) Pool 2×1 Pool 4×4
(3,3)@64, LN, Mish
FC320 Pool 2×1
FC320 GRU 320
GRU 320
FC512, Dropout, FC1 FC512, Dropout FC13 Global pooling
softmax FC64, Tanh, FC13 FC512, Dropout, FC13
softmax
FC13
Timestamp SED Class
Appendix B Impact of the noise decoder on classification
We conducted a detailed analysis to investigate the impact of the noise decoder on classification. Figure 6 presents a class-wise comparison of classification recall performance between the cases with and without the use of the noise decoder. Without the noise decoder, the classification recall for domestic sounds is 49.4%, which is lower than other classes. This is due to the similar time-frequency characteristics of domestic sounds and background noise (from the TAU-SNoise DB [^2] https://zenodo.org/records/6408611), leading the model to misclassify domestic sounds as noise and remove them. However, after adding the noise decoder, the classification recall for domestic sounds increased significantly to 64.9%, showing an improvement of 15.5% points. Furthermore, the recall performance for almost classes improved. These results indicate that although the noise decoder estimates the noise waveform, it significantly influences classification performance.
In Figure 7, the t-SNE analysis further confirmed that without the noise decoder, domestic sounds (represented in purple) have unclear boundaries with other classes, but when the noise decoder is employed, the boundaries between the domestic sound and other classes are more clearly delineated. This suggests that the noise decoder explicitly estimates the noise, allowing the model to better distinguish between foreground sources and noise. In summary, employing a noise decoder addresses the issue in previous approaches where the undenoised background noise degraded classification performance.
Figure 6: Per-class classification performance with and without noise decoder
Figure 7: T-SNE comparison with and without noise decoder
Appendix C Effect of direct and reverb decoder on DoAE
To investigate the effect of estimating direct and reverberant foreground signals separately on DoAE, we conducted a detailed analysis. Estimating direct and reverberant signals separately introduces the dereverberation task in addition to the denoising and separation tasks, which can lead to a decrease in overall separation performance. However, as shown in Figure 8, the LE histogram for DoAE reveals an increase in the number of samples with LE within 10 degrees, while the number of samples with LE greater than 10 degrees decreases. This indicates that even though the direct and reverb decoders are part of the audio decoder, they positively influence DoAE performance.
To understand how the dereverberation task affects DoAE, we performed a cosine similarity analysis between the convolution kernel weights. Figure 9 compares the cosine similarity between the first convolution kernel weights of the DoA decoder and those of the direct and reverb decoders. The results show that the similarity between the DoA decoder and the direct decoder is higher than that between the DoA decoder and the reverb decoder. This suggests that the DoA decoder relies more on the direct components than the reverb components for DoAE. In reverberant environments, DoAE is more challenging compared to anechoic environments. This observation implies that learning to separate direct and reverb components within the object features can improve DoAE by ensuring that the direct components are properly embedded in the features.
Figure 8: Comparison of LE histograms when estimating (a) direct + reverb signals together and (b) direct and reverb separately.
Figure 9: Comparison of cosine similarity between the weight of the first convolution kernel of the DoA decoder and the convolution kernel weights of (a) direct decoder and (b) reverb decoder
Furthermore, to evaluate the performance of dereverberation, we compared the results of reverberant, direct (dereverberation, $\text{SI\text{-}SDR}_{\mathbf{s}}$) and reverb source separation ($\text{SI\text{-}SDR}_{\mathbf{h}}$), as well as the speech-to-reverberation modulation ratio (SRMR) [55]. SRMR is only calculated for classes corresponding to male and female speech. The experimental results demonstrate that the proposed model performs well in dereverberation, achieving performance comparable to full source separation. Additionally, with an SRMR of 6.5, the model indicates strong dereverberation effectiveness.
Table 6: Dereverberation performance of proposed model
SI-SDR $\textbf{SI\text{-}SDR}_{\mathbf{s}}$ $\textbf{SI\text{-}SDR}_{\mathbf{h}}$ SRMR
10.8 10.8 10.1 6.5
Appendix D Detailed analysis on time-varying learnable window
In this section, we present experiments to investigate the role and functionality of the time-varying learnable window. In the audio encoder, the short-time Fourier transform (STFT) is widely used to convert a waveform into a complex spectrogram. Unlike speech separation, which considers only the speech class, USS involves various classes, necessitating dynamic adjustments to the configuration of the window. Although many conventional methods rely on fixed windows [28, 12, 38], this reliance limits their performance [56]. To address this limitation, we propose a time-varying learnable window that can dynamically adjust the window configuration for both USS and ASA.
Table 7: Ablation study results of time-variant learnable window
Variation USS SED DoAE SELD $\downarrow$ Complexities
SI-SDRi $\uparrow$ SDRi $\uparrow$ ER $\downarrow$ F1 $\uparrow$ LE $\downarrow$ LR $\uparrow$ Param. MAC/s
STFT 10.8 11.5 30.0 69.8 19.3 76.0 0.237 8.1 (+96.8) M 99.0 G
1D Conv 7.8 8.8 40.2 59.0 27.1 61.2 0.338 9.0 (+96.8) M 99.2 G
Time-invariant
learnable window
10.7 11.4 31.3 68.1 19.5 75.6 0.246 8.2 (+96.8) M 99.1 G
TDY-CNN 10.5 11.3 29.7 70.0 18.9 76.3 0.235 9.0 (+96.8) M 99.4 G
[…]
Time-variant
learnable window
[…]
Figure 10: Comparison of source separation performance for each class: (a) conventional STFT, (b) time-varying learnable window
We conducted a detailed ablation study on the effect of the time-varying learnable window. The results are presented in Table 7. Initially, we compared encoding methods using complex STFT and learnable 1D convolution kernels. The performance of complex STFT generally exceeded that of the 1D convolution kernel, suggesting that while the 1D convolution kernel extracts learnable features, complex STFT is more effective at explicitly capturing frequency information in the feature representation. Additionally, the time-invariant learnable window learns a fixed window across time, resulting in a window length that is similar to that of the STFT. The time-invariant learnable window can be obtained by performing mean pooling in the temporal dimension before being passed through the linear layers of the time-varying learnable window. Next, we examined TDY-CNN [56], which learns attention weights for each time frame to control how much information is passed through, capturing phoneme variation. However, this approach has the limitation of being unable to adjust the frequency band information. In contrast, the proposed method allows for control over the frequency band by adjusting the width of the main lobe.
Figure 10 presents a comparison of separation performance for each class. The results show a significant improvement in the separation performance of sporadic sounds, such as door and knock events. Despite being transient, the time-varying learnable window effectively captures the characteristics of these sound objects, allowing the model to improve separation performance.
Figure 11: (a) Time-domain waveform of the audio mixture, (b) window length along the time frame, SED results for (c) conventional STFT, (d) time-varying learnable window, (e) ground truth.
Figure 11 shows a comparison of the sample between conventional STFT and the proposed time-varying learnable window. Figure 11(a) shows the time-domain waveform of the audio mixture, and Figure 11(b) plots its frame-wise standard deviation over time. Figure 11(c), (d), and (e) present the SED result of conventional STFT, that of time-varying learnable window, and ground truth, respectively, with binary values between 0 and 1 (black indicates 0, white indicates 1). The results show that with conventional STFT, object 1 was not captured, and object 3 failed in both classification and timestamp estimation. However, using the proposed method, the model can detect object 1, which improves the sound event detection performance of object 3.
Finally, Figure 12 shows a GradCAM [57] analysis of the separated object features immediately after the object splitter, to examine how well the object features encapsulate the properties of the corresponding sound objects. The model with conventional STFT struggled with object separation, as regions of speech are still embedded within telephone regions. In contrast, the proposed method more effectively isolates the telephone regions by suppressing most of the speech-related components. This demonstrates that the proposed approach effectively separates sound objects based on auditory information.
Figure 12: (a) Spectrogram of the audio mixture, GradCAM results for (b) conventional STFT, (c) time-varying learnable window.
Appendix E Detailed study on Chain-of-Inference
Figure 13: SED result (a) without CoI, (b) with CoI, (c) ground truth, and (d) DoAE result of object 1 without CoI model, (e) DoAE of object 3 without CoI, multiclue attention map for (f) DoA query, (g) SED query
In this section, we conduct a case study to understand the underlying mechanism of the CoI. It was observed that the model struggled with SED before applying the CoI, whereas the SED performance improved after its application. We then analyzed the fundamental reasons for this improvement. Figure 13(a), (b), and (c) present the SED output for the model without CoI, the model with CoI, and the ground truth, respectively. By comparing the performance before and after applying CoI, an improvement is observed in the SED estimation for object 1 and object 3. To investigate the cause of this difference, we examined the DoAE results for object 1 and object 3 before applying CoI in Figure 13(d) and (e), and the multi-clue attention maps with SED as the query and DoA as the query in Figure 13(f) and (g), respectively. In the DoAE results, for object 1, DoA vectors below the onset threshold (0.5) suggest the potential presence of an object, while for object 3, the estimated DoA closely matches the ground truth. This indicates that the classification of object 3 was refined through CoI by leveraging the nearly accurate DoAE. Furthermore, examining the attention maps, the attention map with SED as the query identified four objects, whereas the one with DoA as the query identified five. This demonstrates the successful correction of object 1, which was mistakenly classified as silence, by utilizing DoA information.
Appendix F Comparison with SOTA models for ASA2 dataset and real-world demonstration
Comparison USS and SELD performance on ASA2 dataset We compared the performance of the proposed DeepASA model on the ASA2 dataset with that of other SOTA models in Table 8. First, the proposed model achieved superior USS performance compared to other models [26, 46, 48, 38], even though it operates with a MIMO setup and performs a dereverberation task. Despite having a lightweight feature aggregation module compared to DeFT-Mamba, the model still achieved SOTA performance. This improvement is likely due to the ability of the model to utilize various auditory information, such as class, timestamp, DoA, and noise, enabling better separation of foreground sources. Next, in terms of SELD performance, the proposed model significantly outperformed existing models. This is because ASA2 consists entirely of polyphonic audio, and earlier approaches that estimate the outputs directly from mixtures are unsuitable for such overlapped sources. Our model avoids this limitation by separating the objects.
Table 8: Comparison with SOTA models on ASA2 dataset
model USS SED DoAE SELD $\downarrow$ Complexities
SI-SDRi $\uparrow$ SDRi $\uparrow$ ER $\downarrow$ F1 $\uparrow$ LE $\downarrow$ LR $\uparrow$ Param. MAC/s
ByteDance-uss [7] 5.2 8.3 - - - - - 28 (+80.7) M 40.1 G
TF-GridNet [46] 8.7 10.7 - - - - - 14.7 M 462 G
SpatialNet [48] 9.6 10.2 - - - - - 7.3 M 71.8 G
DeFT-Mamba [38] 10.4 11.3 - - - - - 3.6 M 83.8 G
EINV2 [52] - - 48.5 39.5 27.1 51.5 0.431 51.5 M 6.7 G
ResNet Conformer [58] - - 45.7 41.0 26.6 53.7 0.414 13.6 M 7.6 G
SELD-Mamba [54] - - 43.5 42.7 25.5 56.7 0.396 75.1 M 4.3 G
MFF-EINV2 [49] - - 42.1 43.2 25.8 60.7 0.381 54.8 M 14.3 G
[…]
Real-world demonstration To evaluate the applicability of the proposed method and dataset in real-world scenarios, we experimented using a pre-trained model without fine-tuning. The experiment was carried out in a real office environment, where three sound objects (male speech, music, and a telephone) were placed. Among these, the male speech was a moving source. The miniDSP ambiMIK-1 was utilized for capturing the sound, which is different from the microphone configuration for the pre-training dataset. The results demonstrate that the pre-trained DeepASA model can be applied to USS, SED, and DoAE tasks in real-world conditions, confirming its feasibility. A demo video of the real-world experiment and example results from the ASA2 dataset are available on the page linked below [^3] https://huggingface.co/spaces/donghoney22/DeepASA.